In earned value analysis the "schedule variance" is the earned value minus the planned value. Consider the following situation. A project has a planned schedule of 12 months with an estimated project cost of $800,000 . At six months into the project, the project manager quits and you are asked to take over the project. You discover that the accumulated costs are already $200,000 but only 25% of the work is complete. What is the schedule variance for this project?

12 months

$600,000

-$200,000

6 months
Answer #1

Earned value = 25% of work is completed = 25% of 8,00,000 = $2,00,000

Planned value = (completed time/total time planned) * total budget =(6/12) * 8,00,000 = $4,00,000

Schedule variance = Earned value - Planned value = 2,00,000 - 4,00,000 = -$2,00,000

Hence the correct answer is 3rd option i.e., -$200,000